OpenStudy (anonymous):

x-5 1/7=12 8/9 Show Work .

OpenStudy (anonymous):

\[x-(5 + 1/7) = 12 + 8/9\]\[x - (36/7) = 116/9\]\[x =\frac{ 116 }{ 9 } + \frac{ 36 }{ 7 }\]\[x =\frac{ 116 }{ 9 }(\frac{ 7 }{ 7 }) + \frac{ 36 }{ 7 }(\frac{ 9 }{ 9 })\]\[x = \frac{ 324 }{ 63 } + \frac{ 812 }{ 63 }\]\[x = \frac{ 1141 }{ 63 }\]\[x = \frac{ 163 }{ 9 }\]\[x = (18+1/9)\] if that makes sense